log2 (base of 2) of x + 5 (no parentheses) = 8 - log2 (base of 2) of (x+7)
Answer by stanbon(75887)

log2 x + 5 = 8 - log2 (x+7)
log2 x + log2 (x+7) = 3
log2[x(x+7)] = 3
x(x+7) = 2^3
x^2+7x = 8
x^2+7x-8 = 0
(x-1)(x+8)=0
x=1 or x= -8
-----------------------
Checking these answers in the original equation:
x cannot be -8 as log2 (-8) is meaningless.
Checking x=1:
log2 1 + 5 = 8 - log2 (1+8)
0 + 5 = 8 - 3
5 = 5
x=1 is the only solution
=======================
Cheers,
Stan H.